### v3

# Enhancing the Robustness of LLM-Generated Code: Empirical Study and Framework

Abstract.
Ensuring the robustness of code generated by large language models (LLMs) is crucial for real-world reliability. However, existing evaluations predominantly focus on correctness, often neglecting key robustness concerns such as missing input validation and insufficient error handling.
In this paper, we present the first empirical study on the robustness of LLM-generated code using the CoderEval benchmark. We introduce novel robustness metrics and analyze four state-of-the-art code LLMs, revealing that, on average, 43.1% of their generated code is less robust than human-written counterparts. Notably, over 90% of robustness deficiencies stem from missing conditional checks, with 70% of these omissions occurring in the first line of code. Additionally, in 69% of cases where a conditional statement is necessary but absent, the “if” token still ranks third or higher in the model’s predicted token probabilities, indicating an implicit recognition of control structures.
Building on these findings, we propose RobGen, a framework designed to enhance code robustness without requiring model retraining. RobGen leverages two model-agnostic techniques: RobGen-Adj, which dynamically adjusts token probabilities during decoding to encourage the inclusion of control structures, and RobGen-Ins, which improves generated code by inserting missing conditionals after generation. Experimental results demonstrate that RobGen reduces the proportion of less robust model-generated code by 20.0%, significantly enhancing code reliability across diverse tasks. As a lightweight and adaptable solution, RobGen effectively mitigates robustness challenges in LLM-generated code. All code and data are available at https://github.com/SYSUSELab/RobGen.
Code generation, Robustness, Large language models
copyright: acmlicensedjournalyear: 2024
1. Introduction
Automatic code generation, which involves synthesizing code snippets that fulfill specified requirements, has become a crucial aspect of modern software engineering (OpenAI, 2021; Zhang et al., 2023; Radford et al., 2018; Dong et al., 2023; Jiang et al., 2023; Zan et al., 2023). The emergence of large language models for code (Code LLMs), such as CodeLlama (Roziere et al., 2023a), StarCoder (Li et al., 2023), DeekSeekCoder (Guo et al., 2024b) and Qwen2.5-Coder (Yang et al., 2024b), has significantly advanced the capabilities of automated code generation (Chen et al., 2021b; Nijkamp et al., 2022; Roziere et al., 2023b; Li et al., 2023; Guo et al., 2024c). While prior research has primarily focused on improving the correctness of generated code (Ugare et al., 2024; Su and McMillan, 2024; Islam et al., 2024a), robustness remains relatively underexplored. Ensuring robustness is essential for handling edge cases, invalid inputs, and unexpected execution scenarios.
Although many studies measure LLM-generated code correctness using pass@k, such metrics fall short of capturing robustness. Even if a model passes all tests, it may still lack necessary robustness checks (Liu et al., 2024c). Conversely, a model might fail some tests while its core logic is correct, but it may miss critical validations such as handling empty inputs (see Figure10(a)), the code fails to pass the task tests due to missing input parameter checks for “str” and “searchStrArray”. These issues underscore that unit test pass rates alone do not fully reflect code robustness. This gap motivates our study, where we develop new robustness metrics and conduct an in-depth empirical investigation into the robustness deficiencies of LLM-generated code.
Recently, Liu et al. (Liu et al., 2024b) conducted a systematic analysis of ChatGPT-generated code, evaluating its correctness while also identifying potential quality issues. Zhong et al. (Zhong and Wang, 2024) introduced the RobustAPI dataset to assess the reliability and robustness of LLM-generated code. However, these studies did not deeply explore the robustness of model-generated code. To improve code robustness, Zhang et al. (Zhang et al., 2024) proposed Seeker, a multi-agent framework to generate high-quality exception-handling code. However, the use of a multi-agent system requires multiple LLM invocations, leading to significant time overhead.
Empirical Study. This work presents the first comprehensive study on the robustness of LLM-generated code using CoderEval, a benchmark for complex repository-level code generation. New robustness metrics are designed to assess code properties directly, moving beyond reliance on unit tests by incorporating comparisons with human-written ground truth. The study analyzes four Code LLMs, investigating (1) the robustness gap between generated and human-written code, (2) prevalent patterns of robustness issues, (3) the specific locations where these issues arise, and (4) whether LLMs inherently recognize the need for robustness checks.
Our empirical study reveals that on average 43.1% of LLM-generated code is less robust than human-written code, with significant room for improvement. We identify nine distinct patterns of robustness issues, 90% of which are related to missing conditional checks. Further analysis shows that 70% of these issues occur in the first line of the generated code. Additionally, in these patterns, the ”if” token ranks third or higher in 69% of the cases where conditionals should be generated but are missing.
Plug-in Framework. Building on our findings, we introduce RobGen, a plug-in framework for enhancing the robustness of LLM-generated code without retraining. RobGen consists of two model-agnostic independent techniques: RobGen-Adj and RobGen-Ins. RobGen-Adj operates during decoding, dynamically adjusting token logits to encourage the generation of essential control structures. RobGen-Ins refines code post-generation by identifying and inserting missing condition checks. As a lightweight and adaptable solution, RobGen effectively improves code reliability across different models and tasks while fully leveraging the model’s inherent capabilities. We propose the Robustness Relative Index (RRI) to measure the robustness difference between generated and reference code. An RRI below zero (RRI ¡ 0) means the model-generated code is less robust. Experimental results across five models show that RobGen reduces the proportion of code with RRI ¡ 0 by 20%, demonstrating its effectiveness.
We summarize the main contributions of this paper as follows.
  • Conducting the first in-depth empirical analysis of robustness in LLM-generated code.
  • Proposing quantitative metrics to assess robustness, focusing on edge cases, invalid inputs, and error handling.
  • Introducing RobGen, a plug-in framework with RobGen-Adj and RobGen-Ins, two lightweight, model-agnostic techniques that refine control structures during and after generation without retraining.
2. Empirical Study
To assess the robustness of LLM-generated code, we conduct an empirical study with the following research questions (RQs):
  • RQ1: How robust is the code generated by LLMs compared to human-written code? We evaluate LLM-generated code robustness by comparing it to human-written code, focusing on handling boundary conditions, invalid inputs, and unexpected runtime scenarios.
  • RQ2: What are the common patterns of robustness issues in LLM-generated code? We analyze cases where LLM-generated code is less robust than the reference, identifying recurring issues like missing null checks to reveal weaknesses and inform improvement strategies.
  • RQ3: Where do robustness issues tend to occur in LLM-generated code? We investigate the distribution of robustness issues within the generated code at the line level.
  • RQ4: Do LLMs recognize the need for condition statements to improve robustness? We analyze token probability distributions at positions where an if-statement is expected to determine whether LLMs inherently recognize the necessity of condition statements to enhance code robustness.
2.1. Experiment Setup
This section outlines the empirical study settings, including model selection, task selection, and implementation details.
2.1.1. Model Selection
We selected four mainstream and representative open-source Code LLMs that have demonstrated strong performance in code generation: Qwen2.5-Coder-1.5B  and Qwen2.5-Coder-7B  (September 2024) (Hui et al., 2024), as well as DeepSeekCoder-1.3B  and DeepSeekCoder-6.7B  (January 2024) (Guo et al., 2024c). These models are instruction-tuned versions of their base models, enhancing their ability to follow instructions and improving their performance in code generation tasks (Yang et al., 2024b; DeepSeek-AI, 2024).
All models were sourced from official sources and used according to their guidelines. Due to resource constraints, we focus on models with fewer than 7B parameters. Given the limitations of smaller models, ensuring the robustness of their generated code becomes even more critical. Using open-source models also allows for detailed analysis of token probability distributions, offering valuable insights into model behavior.
2.1.2. Task Selection
To evaluate LLM-generated code robustness, we selected CoderEval (Yu et al., 2024) as our benchmark. Unlike OpenAI’s HumanEval (Chen et al., 2021a), which focuses on function-level generation without context, CoderEval offers a more realistic setting with repository-level context, increasing complexity by requiring models to handle dependencies and produce robust code. CoderEval includes 230 Java tasks from real-world open-source projects, each with a docstring, function signature, ground truth implementation, and unit tests. Java’s widespread use in enterprise applications (Ouh et al., 2023), static type system, and strict exception handling (St. Amour and Tilevich, 2024) make it ideal for assessing robustness in LLMs.
2.1.3. Implementation Details
We configure all Code LLMs according to the official guides, using the default settings. To ensure a fair comparison, the maximum token limit is set to 300. During the inference stage, we employ a greedy sampling strategy (Goodfellow et al., 2016) to avoid randomness across the four Code LLMs. All experiments are conducted on a machine equipped with 512 GB of RAM and an Nvidia A800 GPU with 80 GB of memory.
Figure 1 shows an example of the prompts used in our experiments, based on previous work (Guo et al., 2024a). For each task, we combine the task description, context, and method signature as input and ask the LLMs to generate the full method. In some cases, models generate extra code beyond the given method signature, as noted in prior studies (Guo et al., 2024a). To improve performance, we use a rule-based matching approach to filter out unnecessary code.
Figure 1. Code Generation Prompt Used in Experiments
2.2. RQ1: Robustness of LLM-Generated vs. Human Code
To answer this RQ, we compare the robustness of LLM-generated code with the ground truth implementations, which are human-written code from real-world projects.
2.2.1. Analysis Target Selection
As described in Section 2.1, we use 230 Java coding tasks from CoderEval and generate one code snippet per task using different models with greedy decoding. Table 1 presents the Compile@1 (i.e., the number of generated code snippets that successfully compile without errors) and Pass@1 (i.e., the number of snippets that pass all test cases), along with the corresponding counts for each model.
We first filter out LLM-generated code that fail to compile as such code is often incomplete or contains undefined methods and variables. However, our evaluation extends beyond snippets that pass all test cases. We observe that some code failures result not from logic errors but from missing robustness checks, such as improper handling of boundary conditions (e.g., A <= B vs. A < B). These cases, which fail tests due to a lack of robustness, are central to our RQ1 analysis and are crucial for assessing the robustness of model-generated code.
Table 1. Compile@1 and Pass@1 for Studied LLMs
Model Compile@1 Pass@1
DeepSeekCoder-1.3B 0.67 (153) 0.34 (79)
DeepSeekCoder-6.7B 0.75 (172) 0.46 (105)
Qwen2.5-Coder-1.5B 0.67 (155) 0.39 (90)
Qwen2.5-Coder-7B 0.74 (171) 0.49 (112)
2.2.2. Analysis Metrics
To evaluate the robustness of generated code, we adopt principles from defensive programming (McConnell, 2004), which emphasizes boundary checking, exception handling, and input validation to ensure reliable software behavior. Robust code anticipates and mitigates potential errors by systematically handling abnormal conditions. Our analysis focuses on two key aspects: control expressions and exception handling.
Control Expressions Analysis. To evaluate the structure and complexity of control logic, we analyze atomic Boolean expressions, which are the smallest evaluable Boolean expressions found in conditional statements (e.g., if, for, while). These include simple conditions such as x > 5 or isValid() that serve as fundamental decision-making units in code. Using abstract syntax trees (ASTs) extracted via tree-sitter (Hendrik van Antwerpen, 2022), we systematically identify and extract all atomic Boolean expressions from the generated and reference code. This enables us to quantitatively assess the robustness of control structures.
Atomic Boolean Expression Extraction Method. When extracting atomic Boolean expressions, if the expression involves local variables, we replace them with their corresponding values or expressions in the code. Specifically, if a local variable is assigned a value from a method call or an operation, the atomic Boolean expression will be rewritten to reflect that operation directly, avoiding any reference to intermediate variables. For example, consider the code int len = a.length; and if (len > 0). In this case, the atomic Boolean expression extracted from the if statement would be a.length > 0 instead of len > 0. This ensures that the extracted atomic Boolean expressions directly reflect the logic being tested in the code, without unnecessary intermediate variables, making the analysis more accurate.
The following metrics are defined to facilitate this analysis:
Average Atomic Boolean Expressions (AvgABE): The AvgABE measures the average number of atomic Boolean expressions per code snippet. It is calculated as:
(1) $AvgABE=\frac{1}{N}\sum_{i=1}^{N}|U_{i}|$
where $U_{i}$ represents the set of atomic Boolean expressions in code snippet $i$, and $N$ is the total number of snippets. A higher AvgABE value suggests that the code contains more explicit conditions, potentially indicating stronger robustness through enhanced boundary checking and input validation.
Atomic Boolean Expression Similarity (ABES): We define ABES to measure the proportion of shared atomic Boolean expressions between the generated code and the reference code. Given the sets of atomic Boolean expressions in the generated code ($U_{A}$) and the reference code ($U_{B}$), ABES is computed as:
(2) $ABES(A,B)=\frac{|U_{A}\cap U_{B}|}{|U_{B}|}$
When calculating the intersection between $U_{A}$ and $U_{B}$, we account for potential differences in how semantically equivalent atomic Boolean expressions are written. For instance, expressions like len > 0 and len $\geq$ 1, or File.exists() == True and File.exists() $\neq$ False, while syntactically different, may express equivalent logic. To handle these cases, we apply a set of predefined rules to perform approximate matching between these expressions. This metric quantifies the extent to which the generated code aligns with human-written code in terms of control expression structure. A higher ABES value suggests that the generated code adopts similar robustness practices, such as boundary checks and validation logic, as seen in human-written code.
Relative Robustness Index (RRI): The RRI quantifies the relative robustness of generated code $A$ compared to reference code $B$ by analyzing atomic Boolean expressions in control structures. This metric not only measures the extent to which the generated code preserves the robustness mechanisms of the reference but also evaluates whether it introduces additional safeguards. A higher RRI indicates that the generated code enhances robustness by incorporating new checks, while a lower value suggests a loss of essential robustness features.
As defined in Eq 3,
(3) $RRI(A,B)=\frac{|U_{A}\cap U_{B}|+\alpha|U_{\text{extra}}|}{|U_{B}|}-1$
where $U_{A}$ and $U_{B}$ represent the sets of atomic Boolean expressions in the generated and reference code, respectively, and $U_{\text{extra}}=U_{A}\setminus U_{B}$ captures additional robustness checks unique to the generated code. The weighting factor $\alpha$ (set to 1) modulates the impact of these extra expressions. An RRI of 0 indicates that the generated and reference code exhibit comparable robustness. A positive RRI suggests that the generated code strengthens robustness beyond the reference, while a negative value implies that the generated code omits some robustness measures present in the reference. By jointly considering both retained and newly introduced robustness checks, RRI provides a structured and interpretable measure for evaluating robustness improvements or degradations in generated code.
Exception Handling Analysis Exception handling is a key indicator of robustness. Our analysis focuses on two aspects: the adoption rate of exception handling and its consistency with the reference code.
Exception Handling Adoption Rate (EHAR): EHAR quantifies the proportion of code snippets that include try-catch blocks:
(4) $EHAR=\frac{\text{\# of snippets with try-catch}}{\text{Total \# of snippets}}$
A higher EHAR suggests a greater reliance on exception handling, which can be beneficial for capturing runtime errors but may indicate excessive dependence on reactive mechanisms rather than preventive checks.
Exception Handling Consistency (EHC): EHC measures the alignment of exception handling strategies between generated and reference code:
(5) $EHC(A,B)=\frac{|EH_{A}\cap EH_{B}|}{|EH_{B}|}$
where $EH_{A}$ and $EH_{B}$ represent the sets of snippets using try-catch in the generated and reference code, respectively. A lower EHC value suggests a deviation from human-written exception handling patterns, indicating either overuse or underuse of structured error handling.
Table 2. Comparison of Robustness Metrics Across Models: “DSC” for DeepSeekCoder and “QWC” for Qwen2.5-Coder
Model EHAR EHC AvgABE ABSE
Gen GT Gen GT
DSC-1.3B 0.02 0.03 0.40 1.39 2.04 0.42
DSC-6.7B 0.03 0.02 0.75 1.40 2.10 0.46
QWC-1.5B 0.01 0.03 0.25 1.62 2.10 0.48
QWC-7B 0.03 0.03 0.80 1.78 2.02 0.48
Figure 2. RRI of Different LLMs. RRI ¡ 0: Less robust; RRI ¿ 0: More robust (vs. reference).
2.2.3. Results
For the selected LLM-generated code (that can compile), we computed control expression and exception handling metrics relative to the reference code. The results are presented in Table 2 and Figure 2.
Control Expression Analysis. The AvgABE metric quantifies the average number of atomic Boolean expressions per code snippet. As indicated in Table 2, for the ground truth code, the AvgABE values across different subsets of model-generated code consistently remain around 2.0, suggesting that each ground truth snippet contains approximately two atomic Boolean expressions on average. In contrast, model-generated code tends to exhibit lower AvgABE values. Notably, DeepSeekCoder-1.3B demonstrates the lowest AvgABE value of 1.39. Among the evaluated models, Qwen2.5-Coder-7B performs relatively well, achieving an AvgABE of 1.78.
The ABES metric measures the proportion of atomic Boolean expressions shared between the generated and reference code, relative to the total number of atomic Boolean expressions in the reference code. The results indicate that, across different models, ABES values are relatively consistent, suggesting that despite variations in the average number of atomic Boolean expressions across models, the overlap between generated and reference code in this aspect remains stable. The only exception is DeepSeekCoder-1.3B, which exhibits a significantly lower ABES value of 0.42. For other models, the overlap remains around 50%.
Figure 2 shows the distribution of RRI ¡ 0, RRI = 0, and RRI ¿ 0 across models and tasks, representing code with lower robustness, equal robustness, and better robustness compared to the reference code, respectively. Overall, 33.9%-52.3% (43.1% on average) of the code is RRI¡0, 36.0%-49.7% (42.9% on average) has an RRI of 0, and 11.8%-16.4% (14.1% on average) has an RRI ¿ 0. The gap between the proportions of RRI ¡ 0 and RRI ¿ 0 highlights significant potential for improvement.
For Qwen2.5-Coder-1.5B, 45.8% of compiled code is less robust than the reference code, while for Qwen2.5-Coder-7B, this proportion decreases to 33.9%. This suggests that increasing the parameter size of the Qwen2.5-Coder models leads to improved robustness in generated code. These results suggest that, overall, larger models tend to generate more robust code. When comparing models of similar parameter sizes, Qwen2.5-Coder exhibits superior robustness in code generation compared to DeepSeekCoder. Specifically, for Qwen2.5-Coder-1.5B, the proportion of code with RRI = 0 is 38.7%, and the proportion with RRI ¿ 0 is 15.5%. In contrast, for DeepSeekCoder-1.3B, these proportions are 36.0% and 11.8%, respectively, indicating slightly weaker robustness.
Exception Handling Analysis. The EHAR metric quantifies the proportion of code with try-catch blocks. As shown in Table 2, the EHAR values for ground truth and model-generated code remain consistently around 0.03, with minimal variation. It indicates that about 3% of the ground truth code includes explicit exception handling, reflecting the low prevalence of try-catch. This is reasonable, as best practices discourage overusing try-catch, with foreseeable exceptions handled using guard clauses. Similarly, the EHAR values for model-generated code are low. Qwen2.5-Coder-1.5B has a value of 0.01 (1%), while Qwen2.5-Coder-7B rises to 0.03, matching the prevalence in human-written code.
The EHC metric measures the consistency of exception handling between generated and reference code. Qwen2.5-Coder-7B scores 0.8, indicating strong alignment with human-authored patterns, while Qwen2.5-Coder-1.5B scores 0.25, suggesting difficulty in replicating ground truth exception handling.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: Analysis of robustness-related metrics shows that, compared to human-written code, LLM-generated code is suboptimal, with an average of 43.1% being less robust. There is significant potential for improvement. Among various factors, control expressions require more focus than exception handling, as they are more common in real-world code and exhibit a larger performance gap in LLM-generated code.
2.3. RQ2: Patterns of Robustness Issues in LLM-Generated Code
We investigate common patterns of robustness issues in LLM-generated code by analyzing cases where the generated code is less robust than the reference implementation.
2.3.1. Data Selection
The dataset used for this analysis originates from RQ1, consisting of 920 generated code snippets from four models (230 snippets per model). To efficiently identify robustness issues, we leverage the Relative Robustness Index (RRI) from RQ1 2.2.2, selecting only snippets where $RRI<0$ for further analysis. Given the high cost of full manual analysis, we use RRI as an initial filtering step to retain only code snippets exhibiting lower robustness than the reference implementation, resulting in a refined dataset with 437 code snippets for detailed manual inspection.
Further, we manually examine each retained snippet and filter out cases where the generated code is entirely incorrect and incomparable to the reference. For example, a snippet with compilation errors due to formatting inconsistencies in the return value but otherwise correct logic is kept, while one that fails to implement the required functionality is discarded. As a result, 400 code snippets are kept for further analysis.
2.3.2. Analysis Process
For each code snippet, we analyze whether robustness issues exist in the code snippet and what category can it be categorized into. For example, some robustness issues are arise from flaws in Boolean logic expressions affecting control flow, such as missing conditions, incorrect conditions, or inadequate input validation. Using an open coding approach (Strauss and Corbin, 1990), we iteratively categorize these issues, refining the categories as needed. A single snippet may exhibit multiple issues and receive multiple labels. Two authors independently annotate the data, resolving disagreements through discussion and majority consensus.
To ensure accurate assessment, we consider additional contextual information, including reference code, surrounding class context, compilation feedback, and test results. For complex cases where LLM-generated code implements multiple functionalities, each part is individually compared against the reference implementation. The comparison focuses on semantic equivalence rather than surface-level differences; for instance, expressions like “child == null” and “childNode == null” are considered equivalent if they serve the same logical purpose.
Figure 3. Category of Poor Robustness Pattern
2.3.3. Results
Figure 4. Distribution of Robustness Issue Patterns
Based on our observations, we divided the situations where code lacks robustness into two main categories: Missing and Errors, as shown in Figure 3.
  • Missing: This indicates that the model has not recognized the need for an additional check to improve the code’s robustness at this point.
  • Error: An error occurs when the model identifies the necessity of generating a robustness check but produces an incorrect condition, thereby compromising the functionality or even the correct execution of the code.
For the Missing category, we further divided it into seven subcategories:
  • Missing Null Checks: A well-written piece of code should check input variables for null pointers as much as possible to prevent operations on null objects that could crash the system, especially in complex software systems where receiving a null object is common in certain modules.
  • Missing Specific Value Checks: A typical example of this is checking whether an array’s length is zero. In engineering practice, checking the length of an array effectively prevents out-of-bounds errors. Generally, checking for specific values in the code is to detect special situations and handle them accordingly.
  • Missing Range Checks: Range checks are usually used to ensure that the value of some variables or indices do not exceed a certain boundary to avoid errors and exceptions.
  • Missing Boolean Value Checks: Some project code may contain global Boolean variables that indicate the current state of the project. Checking these global Boolean variables helps the program make decisions based on the current state of the project.
  • Missing Type Checks: Type checking directly prevents incorrect operations on types or helps the program execute different actions based on the variable’s type.
  • Missing Assertions: In some code, assertions are used for defensive programming. The role of assertions is similar to that of If statements, and they are often found frequently in unit test code. We observed that the model tends to use If statements for defense when generating non-test unit code, rather than using assertions.
  • Missing Error Handling: The try-catch structure is essential for handling exceptions in I/O operations, file handling, and directory creation, preventing crashes and ensuring proper error management. However, we observed that in critical scenarios requiring try-catch for stability, the model sometimes fails to generate it correctly.
For the Error category, we further divided it into two subcategories:
  • Errorous Expression: CodeLLM may generate code with undefined methods or variables, causing compilation failures. We also observed this issue in boolean atomic expressions, affecting code correctness and reliability.
  • Inconsistent Expression: Expecting the model’s generated code to be identical to standard code is unrealistic. The inconsistency in this context refers to variations in boundary condition judgments within atomic boolean expressions. For example, expressions like A <= B vs. A < B or len < head vs. len < size illustrate such inconsistencies.
We analyzed the distribution of the previously identified patterns, and the results are shown in Figure 4. As illustrated, around 90% of the occurrences across the nine patterns are due to Missing Null Checks, Missing Specific Value Checks, Missing Range Checks, and Missing Boolean Value Checks, all of which involve missing conditional checks. Among these, Missing Null Checks is the most common pattern.
Interestingly, for models with fewer parameters, such as DeepSeekCoder-1.3B, Missing Specific Value Checks as the second most frequently observed pattern. In contrast, for larger models, such as DeepSeekCoder-6.7B, Missing Range Checks becomes the second most common pattern. Other patterns within the Missing category, such as Missing Error Handling, as well as those in the Error category, occur at significantly lower frequencies.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: We identified nine distinct robustness patterns. Statistical analysis of their frequencies revealed that over 90% of these issues are related to missing conditional checks, with Missing Null Checks being the most prevalent.
2.4. RQ3: Distribution of Robustness Issues
In this RQ, we analyze the line-level distribution of robustness issues in generated code.
2.4.1. Design
Building on RQ2, where we identified robustness issues in LLM-generated code along with their corresponding patterns, we further analyze the specific locations where these issues manifest. For code containing multiple robustness issues, we record only the first occurrence, as LLMs generate code token by token, meaning earlier tokens influence subsequent ones. Thus, capturing the initial occurrence provides insight into the root cause of robustness deficiencies.
For issues categorized under the Error patterns, the recorded occurrence corresponds to the exact line containing the erroneous construct, such as an incorrect if condition. For issues categorized under the Missing patterns, determining the precise location requires structural alignment with the reference implementation. In such cases, annotators manually examine the reference code to identify where the missing element should have been placed. The designated location is chosen to ensure minimal deviation in control flow between the generated code and the reference implementation. For example, if the generated code omits an essential input validation check and lacks any guard conditions, and the reference code places this check at the beginning, the robustness issue is recorded at the first line. This approach maintains consistency and ensures an objective comparison.
2.4.2. Results
Figure 5. Line-level Distribution of Robustness Issues
Figure 5 illustrates the positional distribution of robustness issues in the generated code for the four evaluated LLMs. Most issues (70%) occur in the first line, primarily due to missing robustness checks, such as input validation. This aligns with expectations, as critical checks like null pointer validation are typically placed at the beginning of a function or method.
A detailed analysis shows that DeepSeekCoder-1.3B exhibits the highest proportion (75.0%) of robustness issues in the first line, while Qwen2.5-Coder-7B has a lower rate (61.3%). Model size also affects the likelihood of first-line robustness issues. As model size increases, the proportion of missing checks decreases. For instance, DeepSeekCoder-1.3B has a 75.0% issue rate, which drops to 70.0% in DeepSeekCoder-6.7B. Similarly, Qwen2.5-Coder-1.5B has a 69.4% rate, reducing to 61.3% in Qwen2.5-Coder-7B. Additionally, DeepSeekCoder models generate a higher proportion of first-line robustness issues compared to Qwen2.5-Coder models, suggesting that differences in training data or model architecture impact robustness of generated code.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: Most robustness issues occur in the first line of the generated code, primarily due to missing robustness checks.
2.5. RQ4: Condition Generation Potential
We analyze token probabilities at expected “if” statement positions to assess whether LLMs inherently recognize the need for control structures essential for robust code, even if they sometimes omit them during greedy decoding.
2.5.1. Design
As indicated in RQ2, a large fraction (90%) of robustness issues in generated code stem from missing condition statements. In earlier experiments, all LLM-generated code was produced using greedy decoding, which always selects the token with the highest logit. To assess the model’s latent understanding, we analyze the ranking distribution of the “if” token in the model’s logit output at positions where an if-statement is expected. Our assumption is that if the “if” token is ranked highly at these positions, the model possesses the inherent capability to generate a complete condition statement, thereby mitigating robustness issues.
Building on the findings from RQ3, we first identify code lines where missing condition checks lead to robustness issues. For each identified line, we determine the expected insertion point—marked by the first non-whitespace token—and exclude cases where an “if” statement already exists. This filtering yields 250 code snippets for analysis. For each snippet, we capture the model’s logit outputs at the insertion point and retain the top 30 ranked tokens for further examination.
2.5.2. Results
Figure 6. “If” token ranking
Figure 6 shows that at positions where an “if” statement is expected, the “if” token consistently ranks within the top five. For Qwen2.5-Coder-1.5B, when the “if” statement is missing, 50.0% of the cases rank “if” second, 31.7% rank it third, and 10% rank it fourth. Similarly, for DeepSeekCoder-6.7B, 43.6% rank “if” second, 29.0% third, and 9.68% fourth.
These results suggest that although the model often assigns a high probability to the “if” token at critical positions, it occasionally fails to generate it under greedy decoding. This gap indicates that the model inherently recognizes the need for condition statements but that its current sampling strategy may not always capture this potential. Adjusting the sampling strategy—for example, by employing temperature sampling, top-k, or top-p methods—could allow the model to better utilize this latent capability, thereby enhancing overall code robustness.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: The high ranking of the “if” token (usually second or third) at positions lacking conditionals suggests that LLMs inherently recognize essential control structures. Optimizing the sampling strategy could help generate more robust code by adding missing “if” branches.
3. Plug-in Framework for Enhancing Code Robustness
Based on our findings, we propose two plug-in techniques within the RobGen framework to enhance code robustness in LLMs: RobGen-Adj (Section 3.1) and RobGen-Ins (Section 3.2), targeting different phases of the code generation process. Figure 7 provides an overview of the RobGen framework. RobGen-Adj is a plug-in that adjusts token probabilities during decoding to encourage the generation of control structures, while RobGen-Ins is a post-generation plug-in that inserts missing conditionals to enhance robustness. RobGen-Adj and RobGen-Ins can operate independently or be combined as needed. RobGen focuses on addressing robustness issues related to missing conditional checks, which account for more than 90% of occurrences (Finding 2).
Figure 7. Overview of RobGen Framework
Figure 8. RobustCoder Prompt
3.1. RobGen-Adj: Decoding Adjustment
Empirical findings from RQ4 (Section 2.5) show that LLMs inherently recognize essential control structures for robustness, as indicated by the high ranking of the “if” token at key positions. Given that token selection is influenced by logits, we hypothesize that dynamically adjusting these logits can guide the model to generate more robust code, especially through conditional statements like “if” branches.
We propose an intervention mechanism that adjusts the logits of the “if” token, assuming that once the model starts generating an “if” statement, it can complete the logic correctly, enhancing code robustness. During token generation, we determine if logit adjustments are needed based on predefined rules and modify the logits accordingly. RobGen-Adj is implemented using the LogitsProcessor interface provided by Transformer(Wolf et al., 2020).
Rule-based Intervention Identification. Unconditionally modifying the “if” token’s logit value at every generation step may disrupt normal code synthesis. To mitigate excessive intervention, we leverage insights from RQ4 (Finding 4), which indicate that the “if” token typically ranks second or third at positions where a conditional statement is missing. Thus, we selectively adjust the logit value only when the if token appears within the top three ranked logits.
If Token Logit Adjustment. To ensure that the “if” token’s logit value surpasses competing tokens, we employ the equation in Eq. 6 to adjust the logit value for the “if” token.
(6) $\text{logit}=\text{logit}+\Delta\times(\text{rank}-1)$
where $\Delta$ denotes the adjustment factor, and rank represents the position of the “if” token in the logits ranking.
For example, an “if” token can be ranked first by suppressing a competing token (e.g., “byte”) that initially ranks higher with a higher logit value. Note that if the LLM predicts that the “if” token has a low probability of being generated based on the context, the adjustment will not affect the result.
To determine $\Delta$, we analyze RQ4 data by calculating the relative difference between the “if” token’s logit and the highest-ranked token, normalized by their rank difference. The adjustment factor $\Delta$ is set to the 90th-percentile of these observed differences, ensuring it surpasses 90% of the values. The derived $\Delta$ values for different models are 2.10, 2.11, 1.63, and 2.29 for DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, Qwen2.5-Coder-1.5B, and Qwen2.5-Coder-7B, respectively. For models not investigated, $\Delta$ can be set to 1.0.
3.2. RobGen-Ins: Post-Generation Insertion
Based on findings from RQ3 (Section 2.4), we observe that robustness issues are often present in the first line of generated code. To address this, we introduce a post-generation insertion mechanism that ensures missing conditional checks, specifically the inclusion of an “if” statement, are handled. It also follows best practices in defensive programming (McConnell, 2004) by adding necessary checks at the beginning of the method implementation.
Given a code generated by LLMs using a specific prompt, we first verify whether the generated code’s first line includes an appropriate check (i.e., an “if” condition). If this check is absent, we trigger an intervention to insert it. The process proceeds as follows:
Check for Insertion Need. We examine the first line of the generated method implementation. If the line lacks an “if” condition (e.g., a missing robustness check), the system triggers a reconstruction of that line using a “Fill-in-the-Middle” (FIM) technique (Bavarian et al., 2022). If an “if” condition is already present, the process terminates.
Conditional Statement Generation Based on FIM. We construct a FIM prompt consisting of three parts: “prefix,” “middle,” and “suffix.” The model’s task is to fill in the placeholder “middle” section (e.g., ¡FIM_Middle¿) by generating the appropriate “if” statement based on the context provided in the “prefix” and “suffix.” The “prefix” includes the method signature along with relevant task-specific context, while the “suffix” consists of the remaining method body following the first line of the generated code.
To guide the model in generating a valid “if” condition, we pre-insert the “if” keyword at the position where the check should appear. This ensures that the model generates code containing the necessary conditional logic.
Post-Generation Filtering. Once the model generates the “if” statement, we perform a quality check. If multiple conditional blocks (e.g., “if-else”) are generated, we retain only the first complete block, as it is typically the highest quality. Subsequent or incorrect blocks are discarded. The generated “if” statement is evaluated based on predefined rules to ensure its correctness and compatibility with the surrounding code. If it passes the validation, it is merged into the original code. Invalid or incomplete conditions are filtered out, preventing errors or inconsistencies from being introduced into the generated code.
This method ensures that only high-quality, relevant “if” statements are inserted, improving the robustness of the generated code while avoiding noise. Note that different LLMs utilize varying special tokens and concatenation strategies for FIM-based training, their FIM templates differ. We employ model-specific templates as provided in the official documentation.
4. Evaluation
In this section, we focus on the following RQs:
  • RQ5: How do different methods perform across various models, and what are the key differences among them?
  • RQ6: How does the efficiency of different generation methods compare to the default code generation process of the model?
4.1. RQ5: Method Comparison
4.1.1. Design
We conducted experiments using our framework on the four models from our empirical analysis, as well as on StarCoder2-7B. To evaluate the effectiveness of our method, we include the RobustCoder Prompt as a baseline in our experiment. RobustCoder Prompt (RP) involves modifying the prompt to encourage the model to generate more robust code by explicitly guiding it toward incorporating necessary checks and handling potential exceptions. These robustness requirements are embedded into the model’s code-generation template, as illustrated in Figure 8. In all methods, we default to employing greedy sampling, with the token limit set to 300. For different models, the adjustment factor $\Delta$ used during generation varied. Specifically, for the four models from our empirical analysis, we adopted empirically determined values, whereas for the StarCoder2-7B, we set the step size to 1. After each prefix, we append the string “if” to guide the model in generating an if-branch. Specifically, since FIM-based generation may produce additional code, we truncate the FIM-generated results and retain only the first if-branch.
4.1.2. Robustness Metric Results
The metric calculation results, presented in Table 3, demonstrate that RobGen effectively enhances the robustness of generated code across five models. When applying RobGen-Adj+Ins, the AvgABE of model-generated code exceeds that of the reference code. For instance, using RobGen-Adj on DeepSeekCoder-6.7B increases the AvgABE from 1.45 to 2.99, while applying RobGen-Ins to Qwen2.5-Coder-1.5B raises the AvgABE from 1.97 to 2.67—both surpassing the reference code’s AvgABE of 2.03. Likewise, both RobGen-Adj and RobGen-Ins significantly improve ABES values. For example, after applying RobGen-Ins, the ABES value of StarCoder2-7B increases from 0.41 to 0.48, whereas Qwen2.5-Coder-7B’s ABES rises from 0.43 to 0.50 with RobGen-Adj. Among all methods, RobGen-Adj+Ins achieves the best performance, indicating that RobGen-Ins and RobGen-Adj effectively complement each other.
The RRI and Pass@1 distributions in Figure 9 highlight the effectiveness of RobGen. For instance, with RobGen-Adj, the proportion of generated code with RRI ¡ 0 for DeepSeekCoder-1.3B drops from 56.5% to 35.7%. Similarly, for StarCoder2-7B, the proportion of code with RRI ¿ 0 rises from 19.6% to 53.0% using RobGen-Adj+Ins.
In comparison, RobustCoder Prompt (RP) is less effective than RobGen. While RP can enhance the robustness of model-generated code, its impact is inconsistent and, in some cases, even detrimental. For instance, after applying RP to StarCoder2-7B, the ABES value remains unchanged, while for Qwen2.5-Coder-1.5B, the AvgABE value decreases from 1.97 to 1.90. These results suggest that adjusting input prompts alone is insufficient for improving the robustness of model-generated code.
Although RobGen significantly enhances the robustness of generated code, its effect on the Pass@1 metric is not always stable. For example, after applying RobGen-Ins, the Pass@1 score for StarCoder2-7B decreases from 41.7 to 39.1. Similarly, Qwen2.5-Coder-1.5B experiences a decline from 39.13 to 36.96 with RobGen-Adj, and Qwen2.5-Coder-7B drops from 48.7 to 47.0 with RobGen-Adj+Ins. However, RobGen can also improve Pass@1 in certain cases: the Pass@1 of DeepSeekCoder-1.3B increases from 34.4 to 37.4 with RobGen-Adj+Ins, while DeepSeekCoder-6.7B improves from 45.7 to 47.8 with RobGen-Adj. We attribute this variability to differences in how models adapt to each method, leading to inconsistent performance across models.
Table 3. AvgABE and ABSE of Diferent Method: “GS” indicates Greedy Sampling and “PGP” indicates Pre-Generation Prompting
Model GS RP RobGen-Adj RobGen-Ins RobGen-Adj+Ins
AvgABE (Ground Truth AvgABE: 2.03)
DSC-1.3B 1.37 1.69 (+0.32) 2.87 (+1.50) 2.32 (+0.95) 3.32 (+1.95)
DSC-6.7B 1.45 1.72 (+0.27) 2.99 (+1.54) 2.42 (+0.97) 3.34 (+1.89)
QWC-1.5B 1.97 1.90 (-0.07) 5.40 (+3.43) 2.67 (+0.70) 5.65 (+3.68)
QWC-7B 1.76 2.06 (+0.30) 2.93 (+1.17) 2.48 (+0.72) 3.22 (+1.46)
STC-7B 1.83 1.83 (+0.00) 4.27 (+2.44) 2.47 (+0.64) 4.49 (+2.66)
ABSE
DSC-1.3B 0.35 0.38 (+0.03) 0.41 (+0.06) 0.43 (+0.08) 0.44 (+0.09)
DSC-6.7B 0.41 0.45 (+0.04) 0.47 (+0.06) 0.48 (+0.07) 0.49 (+0.09)
QWC-1.5B 0.42 0.43 (+0.01) 0.50 (+0.08) 0.47 (+0.05) 0.50 (+0.08)
QWC-7B 0.43 0.44 (+0.01) 0.50 (+0.07) 0.47 (+0.04) 0.51 (+0.08)
STC-7B 0.41 0.41 (+0.00) 0.48 (+0.07) 0.48 (+0.07) 0.50 (+0.09)
Figure 9. RRI and Pass@1 of Different Methods. RRI ¡ 0: Less robust; RRI ¿ 0: More robust (vs. reference).
4.1.3. Example Demonstration of Method-Generated Code
Figure 10 presents the generated results of DeepSeekCoder-1.3B on task “6367670b1a6d9265ec017a0f” and DeepSeekCoder-6.7B on task “636766fe1a6d9265ec017833” using different methods. In Figure (a), we observe that, compared to the ground truth, the output of Greedy Sampling lacks null pointer and empty array checks. When using RobustCoder Prompt, the model still fails to generate these missing robustness checks. However, when applying RobGen-Adj and RobGen-Ins, the model successfully generates the necessary robustness checks, as highlighted in the red box. This demonstrates that RobGen-Adj and RobGen-Ins are more effective than RP in enhancing code robustness.
In Figure (b), we note that the code generated by Greedy Sampling deviates significantly from the Ground Truth. With RobustCoder Prompt, the generated code aligns more closely with the Ground Truth, as RobustCoder Prompt prioritizes avoiding unnecessary try-catch blocks for exception handling. When applying RobGen-Adj, the model-generated code not only fully aligns with the Ground Truth but also includes an additional robustness check to verify file existence before executing subsequent logic. However, with RobGen-Ins, the model introduces a file existence check but fails to fully align with the Ground Truth. This suggests that, compared to RobGen-Ins, RobGen-Adj offers greater flexibility, as it not only aids in generating robustness checks but also guides the model toward producing more accurate and reliable code.
Figure 10. Examples of Code Generated by Different Methods
4.2. RQ6: Efficiency
Table 4. Runtime of the model on 20 tasks (Min): “DSC” indicates DeepSeekCoder, “QWC” indicates Qwen2.5-Coder and “STC” indicates StarCoder2-7B.
Method DSC-1.3B DSC-6.7B QWC-1.5B QWC-7B STC-7B
Greedy Sampling 1.52 3.03 1.83 1.13 3.01
RobGen-Adj 1.53 (+0.6%) 3.05 (+0.7%) 1.83 (+0.0%) 1.18 (4.4%) 3.02 (0.3%)
RobGen-Ins 2.72 (+78.9%) 4.21 (+38.9%) 2.09 (+14.2%) 1.77 (+56.6%) 3.99 (+32.6%)
RobGen-Adj+Ins 1.99 (+30.9%) 3.89 (+28.3%) 1.96 (+7.1%) 1.38 (+22.1%) 3.30 (+9.6%)
4.2.1. Design
We adopted the experimental setup from RQ5 (Section 4.1) and randomly selected 20 tasks from the dataset. Each of the five models executed these tasks using our methods, and we recorded the execution time for each model under each method. To mitigate potential errors introduced by external factors, each method for each model was executed three times, and the final execution time was computed as the average of these runs.
4.2.2. Results
The execution time results are presented in Table4. Among all methods, RobGen-Ins exhibits the highest time overhead. For instance, after applying RobGen-Ins, DeepSeekCoder-1.3B ’s execution time increased from 1.52 to 2.72, representing a 78.9% increase compared to Greedy Sampling. Similarly, for StarCoder2-7B, execution time increased by 32.6% after using RobGen-Ins. This trade-off is considered acceptable, as RobGen-Ins involves an additional model invocation to generate the missing content. In contrast, RobGen-Adj has minimal impact on execution time. For example, when using RobGen-Adj, DeepSeekCoder-6.7B experienced only a 4.4% increase, while Qwen2.5-Coder-1.5B saw a negligible 0.6% increase. Furthermore, RobGen-Adj+Ins results in lower time overhead compared to using RobGen-Ins alone. For instance, when DeepSeekCoder-1.3B employed RobGen-Adj+Ins, the additional time overhead was reduced from 78.9% to 30.9%, demonstrating that RobGen-Adj effectively preemptively addresses robustness issues, thereby reducing the need for additional insertions during RobGen-Ins.
5. Threats to validity
Internally, a potential threat is the use of LLMs in our research. To mitigate biases, we use officially released models, deploy them per publishers’ guidelines, and validate outputs through tests. We also ensure fairness by using consistent prompts and parameters. Another concern is the representativeness and quality of code data. To address this, we use the CoderEval benchmark for realistic code generation. Due to resource constraints, we evaluated only four models and focused on Java, limiting generalizability. More extensive studies are needed. While RobGen was designed based on findings from these four models, we included Starcoder in experiments to demonstrate broader applicability. Externally, we minimize human bias in robustness analysis by using quantifiable metrics instead of manual judgment.
6. Related Work
6.1. LLM-based Code Generation
LLMs have shown remarkable performance in code generation, with applications in automated code generation (Yuan et al., 2023; Ugare et al., 2024; Su and McMillan, 2024; Liu et al., 2024a; Wang et al., 2024c), code translation (Yan et al., 2023; Yang et al., 2024a; Ou et al., 2025; Wang et al., 2024b), commit message generation (Tao et al., 2021; He et al., 2023), unit test generation (Schäfer et al., 2024; Lops et al., 2024; Yuan et al., 2024), and defect localization (An et al., 2024; Ji et al., 2024; Du et al., 2024b). Code-specific LLMs, like DeepSeekCoder (Guo et al., 2024c), StarCoder2 (Lozhkov et al., 2024), and Qwen2.5Coder (Yang et al., 2024b), excel in code-related tasks, with high-quality benchmarks such as HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), Classeval (Du et al., 2024a) and CodeEval (Yu et al., 2023) assessing their capabilities. While most studies focus on improving code generation and evaluating quality with the Pass@k metric (Chen et al., 2021b), some explore code security (Li et al., 2024; Islam et al., 2024b), robustness(Sarker et al., 2024; Awal et al., 2024), hallucination(Zhang et al., 2025) and trustworthy(Wang et al., 2024a). However, existing benchmarks fall short in assessing the robustness of model-generated code. This work compares the robustness of LLM-generated code with human-written code, identifying common deficiencies and offering insights for future improvements.
6.2. Robustness of LLM-generated Code
Earlier research on LLMs focused primarily on accuracy in small-scale tasks, often overlooking code robustness in real-world development. Recently, studies like Liu et al.(Liu et al., 2024b) have analyzed ChatGPT-generated code for correctness and quality issues, while Zhong et al.(Zhong and Wang, 2024) introduced RobustAPI to assess the robustness of LLM-generated code for Java API misuse from Stack Overflow questions. Evaluation results on GPT-4 revealed that 62% of the generated code still exhibited API misuse issues (Zhong and Wang, 2024). Recently, Zhang et al. (Zhang et al., 2024) introduced Seeker, a multi-agent framework for improving exception-handling and robustness through automated detection and optimization with a high cost though.
Previous work lacks in-depth analysis of robustness issues in LLM-generated code, particularly in real-world tasks. Many methods are either impractical or involve high overhead. In contrast, our study provides a deeper empirical analysis and identifies common robustness issues across multiple models. We propose RobGen, a model-agnostic plug-in framework that improves robustness with minimal overhead, enhancing code quality during both decoding and post-generation in a lightweight, adaptable manner.
7. CONCLUSION
In conclusion, this study emphasizes the need for robustness in LLM-generated code, which is often overlooked in favor of correctness. Using the CoderEval benchmark, we identify critical gaps, such as missing null checks. To address these, we propose RobGen, a plug-in framework that enhances code robustness without retraining, through techniques like RobGen-Adj and RobGen-Ins. Our experiments show a 20% reduction in less robust code, demonstrating the effectiveness of this approach. RobGen offers a flexible, model-agnostic solution for improving LLM-generated code reliability across various tasks.
Acknowledgements.